# Sharing Key Semantics in Transformer Makes Efficient Image Restoration

### v1

@@ Abstract
Image Restoration (IR), a classic low-level vision task, has witnessed significant advancements through deep models that effectively model global information. Notably, the Vision Transformers (ViTs) emergence has further propelled these advancements. When computing, the self-attention mechanism, a cornerstone of ViTs, tends to encompass all global cues, even those from semantically unrelated objects or regions. This inclusivity introduces computational inefficiencies, particularly noticeable with high input resolution, as it requires processing irrelevant information, thereby impeding efficiency. Additionally, for IR, it is commonly noted that small segments of a degraded image, particularly those closely aligned semantically, provide particularly relevant information to aid in the restoration process, as they contribute essential contextual cues crucial for accurate reconstruction. To address these challenges, we propose boosting IR’s performance by sharing the key semantics via Transformer for IR (i.e., SemanIR) in this paper. Specifically, SemanIR initially constructs a sparse yet comprehensive key-semantic dictionary within each transformer stage by establishing essential semantic connections for every degraded patch. Subsequently, this dictionary is shared across all subsequent transformer blocks within the same stage. This strategy optimizes attention calculation within each block by focusing exclusively on semantically related components stored in the key-semantic dictionary. As a result, attention calculation achieves linear computational complexity within each window. Extensive experiments across 6 IR tasks confirm the proposed SemanIR’s state-of-the-art performance, quantitatively and qualitatively showcasing advancements.
@@ 1 Introduction
Image restoration (IR) stands as a fundamental task within low-level computer vision, aiming to enhance the quality of images affected by numerous factors, including noise, blur, low resolution, compression artifacts, mosaic patterns, adverse weather conditions, and other forms of distortion. This capability holds broad utility across various domains, facilitating information recovery in medical imaging, surveillance, and satellite imagery. Furthermore, it bolsters downstream vision tasks like object detection, recognition, and tracking [79, 65]. Despite notable progress in recent years, prevalent IR methods encounter challenges in effectively addressing complex distortions or preserving/recovering crucial image details [50]. Achieving high-quality image recovery necessitates meticulous exploration of the rich information present in degraded counterparts.
@@
In this paper, we introduce a novel approach, SemanIR, to address the limitations above. Specifically, within each transformer stage, we first construct a key-semantic dictionary, which stores only the top-k semantically related relations for each given degraded patch with the k-nearest neighbors (i.e., KNN) algorithm. Then, the attention operation within each transformer layer occurs only among the top-k patches. This design brings two main advantages, i.e., 1) Each degraded patch benefits from its semantically similar patches, typically containing comparable contextual or textural information, while excluding the side effects from other patches that contain entirely unrelated information. 2) Compared to the conventional window-wise attention, which built a dense connection between all the patches (Fig. 1(c)) that leads to highly computationally demanding, or a sparse but position-fixed manner (Fig. 1(d)) which introduces irrelevant semantics. Our key-semantic connection (Fig. 1(e)) leads to a sparse yet more meaningful attention operation, which allows our method to achieve the same receptive field as previous ViTs-based methods while maintaining lower computational costs. This is not like previous token merging or pruning methods [6, 95, 74] that may merge unrelated information or prune some semantically related information. In addition, to make the proposed method more efficient, instead of creating a key-semantic dictionary for each transformer layer, we create it just once at the beginning of each transformer stage and then share it with all the following transformer layers within the same stage. This not only largely reduced the computation burden but also made our methods different from other token pruning and merging methods [110, 93], which include dynamic patch skipping/selection within each attention layer or involve an additional offset generation network. In addition, such a sharing strategy allows each degraded patch can be continuously optimized by its semantically related patches within each stage.
It is also worth noting that the implementation of the attention layer of our SemanIR is achieved in three interesting manners (i.e., Triton [20] [^1] Open-source GPU programming tool https://openai.com/research/triton., torch-mask, and torch-gather), which are discussed in our ablation studies. The contributions of this work are:
@@ 2 Related Work
Image Restoration (IR), as a long-standing ill-posed inverse problem, is designed to reconstruct the high-quality image from the corresponding degraded counterpart with numerous applications [77, 3, 51]. Initially, IR was addressed through model-based solutions, involving the search for solutions to specific formulations. However, learning-based approaches have gained much attention with the significant advancements in deep neural networks. Numerous approaches have been developed, including regression-based [53, 43, 52, 15, 50, 108] and generative model-based pipelines [26, 91, 60, 96]. In this paper, we propose a regression-based method for image restoration.
@@
Non-Local Priors Modeling in IR. Tradition model-based IR methods reconstruct the image by regularizing the results (e.g., Tikhonov regularization [28]) with formulaic prior knowledge of natural image distribution. However, it is challenging for these methods to recover realistic detailed results with hand-designed priors. Besides, some other classic method finds that self-similarity is an effective prior, which leads to an impressive performance [7, 18]. Apart from traditional methods, the non-local prior has also been utilized in modern deep learning networks [55, 90, 50, 112], typically captured by the self-attention mechanism. More recently, the overwhelming success of transformers [89] in the natural language processing domain [37] and the classic vision community [21, 8, 86, 94, 10] has led to the development of numerous ViT-based IR methods. These methods aim to enhance the learning ability for modeling non-local priors [52, 97, 50, 14, 16, 102] and consequently archives better performance. Meanwhile, this raises a question: are all non-local priors essential for IR?
@@
Key-Semantic Non-local Prior Exploration for IR. To answer the question, we found many methods demonstrating the effectiveness of modeling key semantics within ViTs. For example, KiT [45] proposed increasing the non-local connectivity between patches at different positions through KNN matching. This approach aims to better capture the non-local relations between the base patch and other patches in each attention calculation. However, it results in significant additional computational costs due to the KNN matching. DRSformer [13] proposed a top-k selection strategy that chooses the most relevant tokens to model the non-local priors for draining after each self-attention calculation without reducing the computation complexity. Similar conclusions can be also drawn from the graph perspective solutions [29, 78, 66, 31, 54, 34] for various IR tasks, like facial expression restoration [58], image denoising [82], and artifact reduction [66]. [34] construct the graph with transformer-based architecture where each patch is connected to all other patches. All these methods suggest that if the semantically related information can be addressed, the degraded image can be restored with better performance. However, the efficiency issue, which is extremely unignorable, remains untouched within the aforementioned methods. It is particularly crucial for ViTs-based image restoration methods, which often need to address high-resolution degraded input images. In this paper, we propose to share key semantics within each transformer stage, an approach validated to be efficient and effective via experimental and theoretical analysis.
@@ 3.3 Discussion
Fixed top-k vs. Random top-k Training Strategies. In the fixed top-k approach, $k$ remains constant at 512 during training. In contrast, in the random top-k method, $k$ is randomly selected from the set $[64,128,192,256,384,512]$. It is important to note that even in the random top-k setting, a fixed k value is maintained for all patches/pixels in each iteration. During inference, $k$ is adjusted based on the computational constraints of users.
@@ 4 Experiments
4.1 Ablation Study
The impact of the implementation of Key-Graph Attention is assessed in terms of (i) Triton, (ii) Torch-Gather, and (iii) Torch-Mask under different numbers of N (various from 512 to 8192) and K (various from 32 to 512). The results of the GPU memory footprint are shown in Tab. 2, which indicate that Torch-Gather brings no redundant computation while requiring a large memory footprint. Though Torch-Mask brings the GPU memory increase, the increment is affordable compared to Torch-Gather and also easy to implement. Triton largely saves the GPU memory while at the cost of slow inference and difficult implementation for the back-propagation process. To optimize the efficiency of our SemanIR, we recommend employing Torch-Mask during training and Triton during inference, striking a balance between the efficiency and the GPU memory requirement.
The Impact of the $k$ in Key-Semantic Dictionary Construction. Three interesting phenomena are observed from the results shown in Fig. 3 regarding the two top-k training strategies (Sec. 3.3). (1) The PSNR can largely increase with the increase of $k$ in a fixed manner. (2) When $k$ reaches a certain number (i.e., 384), the performance improvements become marginal, supporting our statement that only the most semantically related patches contribute significantly to the restoration. (3) The randomly sampled strategy has a very stable and better performance compared to the fixed top-k manner especially when the inference $k$ is fixed to a small number (i.e., 64, 128, 256). We conclude that a random sampled strategy is more general and stable. It can also make the inference process more flexible regarding different computation resources. More ablation results can be found in our Appx. D about the effect of the noise level and quality factor for denoising and JPEG CAR.
@@ 4.1 Ablation Study
Efficiency Analysis. We compare our SemanIR with 4 recent promising methods both the x4 SR and grayscale denoising on the Urban100. The trainable parameters, the FLOPs, the runtime, and the PSNR are reported in Tab. 3. It shows that: (1) Among all the methods, HAT and SemanIR achieve 1st-class PSNR performances, reaching 28.37dB and 28.34dB, while SemanIR-S is much faster and has 41.7% fewer parameters than HAT. (2) SwinIR runs a bit faster than SemanIR-S but with a PSNR loss of 0.89 dB. Compared with CAT, and HAT, our SemanIR-S is faster and more accurate. In addition, the parameters and the FLOPs comparison between SwinIR and our SemanIR for denoising reveals that despite SwinIR having fewer training parameters, its FLOPs significantly surpass ours.
The Impact of One Model is Trained to Handle Multiple Degradation Levels. The experiments were conducted for both denoising and JPEG CAR with both color and grayscale images. For denoising, $\sigma$ is set to $[15,25,50,75]$. For JEPG CAR, QF is set to $[10,20,30,40,50,60,70,80,90]$. The results in Fig. 4 indicate that the PSNR for both tasks across all the datasets, under both color and grayscale settings, decreases when the degraded level increases. It means to maintain a decent generalization ability with one model to handle multiple degradation levels is not easy. Meanwhile, the proposed SemanIR can still outperform other methods on various tasks (See Tab. 7 and Tab. 4), which means that capturing the key semantics is also essential for model’s generalization ability.
4.2 Evaluation of SemanIR on Various IR Tasks
Evaluation on Image deblurring. Tab. 2 shows the quantitative results for single image motion deblurring on synthetic datasets (GoPro [67], HIDE [81]). Compared to the previous state-of-the-art Restormer [97], our SemanIR achieves significant PSNR improvement (i.e., 0.52 dB) on the GoPro dataset and the second-best on the HIDE dataset. Visual results are shown in the Appx. E.2.
Table 7: Color and grayscale image denoising PSNR results.
Method # P Color Grayscale
CBSD68 McMaster Urban100 Set12 BSD68 Urban100
Evaluation on JPEG CAR. The experiments for both grayscale and color images are conducted with 4 image quality factors ranging from 10 to 40 under two settings (i.e., a single model is trained to handle multiple quality factors, and each model for each quality). The quantitative results shown in Tab. 4 and the Appx. E.1 show that our SemanIR achieves the best results on all the test sets cross various quality factors among all the comparison methods for both the color and grayscale images. The visual comparisons in the Appx. E.2 further supports the effectiveness of our method.
Evaluation on Image Denoising. We show color and grayscale image denoising results in Tab. 7 under two settings (i.e., one model for all noise levels $\sigma=\{15,25,50\}$ and each model for each noise level). For a fair comparison, both parameters and accuracy are reported for all the methods. For , our SemanIR performs better on all test sets for color and grayscale image denoising than others. It is worth noting that we outperform DRUNet and Restormer with lower trainable parameters. For another setting, the proposed SemanIR also archives better results on CBSD68 and Urban100 for color image denoising, and on Set12 and Urban100 for grayscale denoising. These interesting comparisons validate the effectiveness of the proposed SemanIR and also indicate that our method has a higher generalization ability. The visual results in Appx. E.2 also support that the proposed SemanIR can remove heavy noise corruption and preserve high-frequency image details, resulting in sharper edges and more natural textures without over-smoothness or over-sharpness problems.
@@ 4.2 Evaluation of SemanIR on Various IR Tasks
Evaluation in AWC. We validate SemanIR in adverse weather conditions, including rain+fog (Test1), snow (SnowTest100K), and raindrops (RainDrop). PSNR is reported in Tab. 6. Our method achieves the best performance on Test1 (i.e., 5.76% improvement) and SnowTest100k-L (i.e. 8.01% improvement), while the second-best PSNR on RainDrop compared to all other methods. See Appx. E.2 for Visual comparisons.
Evaluation on Image Demosaicking. The quantitative results shown in 6 indicate that the proposed SemanIR performs best on both the Kodak and MaMaster test sets, especially 0.05dB and 0.45dB absolute improvement compared to the current state-of-the-art.
Evaluation on SR. For the classical image SR, we compared our SemanIR with both recent lightweight and accurate SR models, and the quantitative results are shown in Tab. 8. Compared to EDT, SemanIR-base achieves significant improvements on Urban100 (i.e., 0.72 dB and 0.76dB for 2$\times$ and 4$\times$ SR) and Manga109 datasets (i.e., 0.22dB and 0.17 dB for 2$\times$ and 4$\times$ SR). Even the SemanIR-small consistently ranks as the runner-up across the majority of test datasets, all while maintaining a reduced number of trainable parameters. Visual results in Fig. 5 and Appx. E.2 also validate the effectiveness of the proposed SemanIR in restoring more details and structural content.

### v2

@@ Abstract
Image Restoration (IR), a classic low-level vision task, has witnessed significant advancements through deep models that effectively model global information. Notably, the emergence of Vision Transformers (ViTs) has further propelled these advancements. When computing, the self-attention mechanism, a cornerstone of ViTs, tends to encompass all global cues, even those from semantically unrelated objects or regions. This inclusivity introduces computational inefficiencies, particularly noticeable with high input resolution, as it requires processing irrelevant information, thereby impeding efficiency. Additionally, for IR, it is commonly noted that small segments of a degraded image, particularly those closely aligned semantically, provide particularly relevant information to aid in the restoration process, as they contribute essential contextual cues crucial for accurate reconstruction. To address these challenges, we propose boosting IR’s performance by sharing the key semantics via Transformer for IR (i.e., SemanIR) in this paper. Specifically, SemanIR initially constructs a sparse yet comprehensive key-semantic dictionary within each transformer stage by establishing essential semantic connections for every degraded patch. Subsequently, this dictionary is shared across all subsequent transformer blocks within the same stage. This strategy optimizes attention calculation within each block by focusing exclusively on semantically related components stored in the key-semantic dictionary. As a result, attention calculation achieves linear computational complexity within each window. Extensive experiments across 6 IR tasks confirm the proposed SemanIR’s state-of-the-art performance, quantitatively and qualitatively showcasing advancements. The visual results, code, and trained models are available at  https://github.com/Amazingren/SemanIR.
@@ 1 Introduction
Image restoration (IR) stands as a fundamental task within low-level computer vision, aiming to enhance the quality of images affected by numerous factors, including noise, blur, low resolution, compression artifacts, mosaic patterns, adverse weather conditions, and other forms of distortion. This capability holds broad utility across various domains, facilitating information recovery in medical imaging, surveillance, and satellite imagery. Furthermore, it bolsters downstream vision tasks like object detection, recognition, and tracking [75, 60, 70]. Despite notable progress in recent years, prevalent IR methods encounter challenges in effectively addressing complex distortions or preserving/recovering crucial image details [48]. Achieving high-quality image recovery necessitates meticulous exploration of the rich information present in degraded counterparts.
@@
In this paper, we introduce a novel approach, SemanIR, to address the limitations above. Specifically, within each transformer stage, we first construct a key-semantic dictionary, which stores only the top-k semantically related relations for each given degraded patch with the k-nearest neighbors (i.e., KNN) algorithm. Then, the attention operation within each transformer layer occurs only among the top-k patches. This design brings two main advantages, i.e., 1) Each degraded patch benefits from its semantically similar patches, typically containing comparable contextual or textural information, while excluding the side effects from other patches that contain entirely unrelated information. 2) Compared to the conventional window-wise attention, which built a dense connection between all the patches (Fig. 1(c)) that leads to highly computationally demanding, or a sparse but position-fixed manner (Fig. 1(d)) which introduces irrelevant semantics. Our key-semantic connection (Fig. 1(e)) leads to a sparse yet more meaningful attention operation, which allows our method to achieve the same receptive field as previous ViTs-based methods while maintaining lower computational costs. This is not like previous token merging or pruning methods [6, 92, 69] that may merge unrelated information or prune some semantically related information. In addition, to make the proposed method more efficient, instead of creating a key-semantic dictionary for each transformer layer, we create it just once at the beginning of each transformer stage and then share it with all the following transformer layers within the same stage. This not only largely reduced the computation burden but also made our methods different from other token pruning and merging methods [107, 90, 108], which include dynamic patch skipping/selection within each attention layer or involve an additional offset generation network. Meanwhile, merging or pruning tokens will lead to a loss of information in corresponding patches, which is not preferred in image restoration [86]. In addition, such a sharing strategy allows each degraded patch to be continuously optimized by its semantically related patches within each stage.
It is also worth noting that the implementation of the attention layer of our SemanIR is achieved in three interesting manners (i.e., Triton [19] [^1] Open-source GPU programming tool https://openai.com/research/triton., torch-mask, and torch-gather), which are discussed in our ablation studies. Overall, our method’s suitability for image restoration comes from the utilization of semantic information, the preservation of the details, and the effective KNN strategy. The contributions of this work are:
@@ 2 Related Work
Image Restoration (IR), as a long-standing ill-posed inverse problem, is designed to reconstruct the high-quality image from the corresponding degraded counterpart with numerous applications [73, 3, 49]. Initially, IR was addressed through model-based solutions, involving the search for solutions to specific formulations. However, learning-based approaches have gained much attention with the significant advancements in deep neural networks. Numerous approaches have been developed, including regression-based [51, 41, 50, 48, 105] and generative model-based pipelines [25, 88, 56, 93]. In this paper, we propose a regression-based method for image restoration.
@@
Non-Local Priors Modeling in IR. Tradition model-based IR methods reconstruct the image by regularizing the results (e.g., Tikhonov regularization [27]) with formulaic prior knowledge of natural image distribution. However, it is challenging for these methods to recover realistic detailed results with hand-designed priors. Besides, some other classic method finds that self-similarity is an effective prior, which leads to an impressive performance [7, 17]. Apart from traditional methods, the non-local prior has also been utilized in modern deep learning networks [87, 48, 110], typically captured by the self-attention mechanism. More recently, the overwhelming success of transformers [85] in the natural language processing domain [35] and the classic vision community [20, 8, 82, 91, 10] has led to the development of numerous ViT-based IR methods. These methods aim to enhance the learning ability for modeling non-local priors [50, 94, 48, 14, 15, 99] and consequently archives better performance. Meanwhile, this raises a question: are all non-local priors essential for IR?
@@
Key-Semantic Non-local Prior Exploration for IR. To answer the question, we found many methods demonstrating the effectiveness of modeling key semantics within ViTs. For example, KiT [43] proposed increasing the non-local connectivity between patches at different positions through KNN matching. This approach aims to better capture the non-local relations between the base patch and other patches in each attention calculation. However, it results in significant additional computational costs due to the KNN matching. DRSformer [13] proposed a top-k selection strategy that chooses the most relevant tokens to model the non-local priors for draining after each self-attention calculation without reducing the computation complexity, since after each attention calculation, the DRSFormer utilized (mask, top-k, scatter) operations at each transformer layer. Consequently, this inevitably increases the computation cost. Similar conclusions can be also drawn from the graph perspective solutions [28, 74, 61, 32] for various IR tasks, like facial expression restoration [54], image denoising [78], and artifact reduction [61]. [32] construct the graph with transformer-based architecture where each patch is connected to all other patches. All these methods suggest that if the semantically related information can be addressed, the degraded image can be restored with better performance. However, the efficiency issue, which is extremely unignorable, remains untouched within the aforementioned methods. It is particularly crucial for ViTs-based image restoration methods, which often need to address high-resolution degraded input images. LaViT [108] reduces computational costs by storing attention scores from a few initial layers and reusing them in subsequent layers. However, this approach does not change the computation cost of attention itself; it merely reuses previously computed scores. In this paper, we propose sharing key semantics within each transformer stage, demonstrating its efficiency and effectiveness through experimental and theoretical analysis. Our method, SemanIR, reduces computation in both training and inference by using a semantic dictionary to filter out irrelevant patches during training and optimizing attention operations with Triton kernels during inference.
@@ 3.3 Discussion
Fixed top-k vs. Random top-k Training Strategies. In the fixed top-k approach, $k$ remains constant at 512 during training. In contrast, in the random top-k method, $k$ is randomly selected from the set $[64,128,192,256,384,512]$. It is important to note that even in the random top-k setting, a fixed k value is maintained for all patches/pixels in each iteration. During inference, the random top-k strategy offers more flexibility and requires training only a single model, making it more user-friendly and less resource-intensive.
@@ 4 Experiments
4.1 Ablation Study
The impact of the implementation of SemanIR Attention is assessed in terms of (i) Triton, (ii) Torch-Gather, and (iii) Torch-Mask under different numbers of N (various from 512 to 8192) and K (various from 32 to 512). The results of the GPU memory footprint are shown in Tab. 2, which indicate that Torch-Gather brings no redundant computation while requiring a large memory footprint. Though Torch-Mask brings the GPU memory increase, the increment is affordable compared to Torch-Gather and also easy to implement. Triton largely saves the GPU memory while at the cost of slow inference and difficult implementation for the back-propagation process. To optimize the efficiency of our SemanIR, we recommend employing Torch-Mask during training and Triton during inference, striking a balance between the efficiency and the GPU memory requirement.
Figure 4: The impact of $k$ with different inference $k$ value.
@@ 4.1 Ablation Study
The Impact of the $k$ in Key-Semantic Dictionary Construction. Three interesting phenomena are observed from the results shown in Fig. 3 regarding the two top-k training strategies (Sec. 3.3). (1) The PSNR can largely increase with the increase of $k$ in a fixed manner. (2) When $k$ reaches a certain number (i.e., 384), the performance improvements become marginal, supporting our statement that only the most semantically related patches contribute significantly to the restoration. (3) The randomly sampled strategy has a very stable and better performance compared to the fixed top-k manner especially when the inference $k$ is fixed to a small number (i.e., 64, 128, 256). We conclude that a random sampled strategy is more general and stable. It can also make the inference process more flexible regarding different computation resources. Meanwhile, we set a query region in the input and provided a detailed comparison from the attention-based activation map together with the input query region in Fig. 4. Fig. 4(a) shows the query region input. Fig. 4(b) displays the activation map generated using standard attention mechanisms. Fig. 4(c-f) illustrate activation maps using our key-semantic dictionary with different $k$ values ([8, 16, 64, 256]) during inference. The comparisons indicate that increasing $k$ allows for connections to more semantically related regions. However, when $k$ is set too high (e.g., $k$ = 256 as shown in Fig. 4(f)), the activation map may include some semantically unrelated regions. This aligns with the findings and the results depicted in Fig 3, where increasing the $k$ beyond a certain point (e.g., from 396 to 512) does not further improve PSNR. More ablation results can be found in our Appx. D about the effect of the noise level and quality factor for denoising and JPEG CAR.
Efficiency Analysis. We compare our SemanIR method with four recent promising approaches on both 4x super-resolution (SR) and grayscale denoising tasks using the Urban100 dataset. Key metrics, including trainable parameters, FLOPs, runtime, and PSNR, are reported in Tab. 3. The results show that HAT and SemanIR achieve top-tier PSNR performance, reaching 28.37 dB and 28.34 dB, respectively, while SemanIR-S is much faster and uses 41.7% fewer parameters than HAT, making it more computationally efficient. SwinIR runs slightly faster than SemanIR-S but at the cost of a significant 0.89 dB loss in PSNR. In comparison, SemanIR-S offers better accuracy and speed than both CAT and HAT, solidifying its effectiveness. Furthermore, the comparison between SwinIR and SemanIR for denoising reveals that, although SwinIR has fewer trainable parameters, its FLOPs are substantially higher, indicating that SemanIR is more efficient in terms of computation. These results highlight that SemanIR-S strikes an optimal balance between performance and efficiency, making it highly competitive for both SR and denoising tasks.
@@
Table 5: Gray image denoising PSNR.
Window Size Set12 BSD68 Urban100
8 31.01 29.49 31.33
16 31.06 29.51 31.45
32 31.17 29.50 31.88
Table 6: Color image denoising PSNR.
Window Size Mcmaster CBSD68 Kodak24 Urban100
8 33.20 31.72 32.84 32.89
16 33.24 31.74 32.86 32.95
32 33.38 31.75 32.97 33.27
The Impact of One Model is Trained to Handle Multiple Degradation Levels. The experiments were conducted for both denoising and JPEG CAR with both color and grayscale images. For denoising, $\sigma$ is set to $[15,25,50,75]$. For JEPG CAR, QF is set to $[10,20,30,40,50,60,70,80,90]$. The results in Fig. 5 indicate that the PSNR for both tasks across all the datasets, under both color and grayscale settings, decreases when the degraded level increases. It means that maintaining a decent generalization ability with one model to handle multiple degradation levels is not easy. Meanwhile, the proposed SemanIR can still outperform other methods on various tasks (See Tab. 7 and Tab. 4), which means that capturing the key semantics is also essential for model’s generalization ability.
The Impact of the window size. The windows indeed contain mixed information from different semantic parts. Yet, it is precisely this semantic distinction that motivates us to develop a selection mechanism for semantic information using KNN. We have conducted ablation studies of the window size (i.e., on both gray and color image denoising with $\sigma=25$). The results are summarized in Tab. 6 and Tab. 6. With the increase of the window size, the semantic relevant information for each token is increased, thus leading to a PSNR gain for different IR tasks. For gray image denoising, larger window sizes lead to improved PSNR performance, consistently yielding higher PSNR values across different datasets. Similarly, in color image denoising, larger window sizes also result in better outcomes. For example, in the McMaster dataset, the PSNR increases from 33.20 dB with a window size of 8 to 33.38 dB with a window size of 32. These results suggest that larger window sizes enhance performance by capturing more contextual information.
4.2 Evaluation of SemanIR on Various IR Tasks
Evaluation on Image deblurring. Tab. 2 shows the quantitative results for single image motion deblurring on synthetic datasets (GoPro [62], HIDE [77]). Compared to the previous state-of-the-art Restormer [94], our SemanIR achieves significant PSNR improvement (i.e., 0.52 dB) on the GoPro dataset and the second-best on the HIDE dataset. The visual results are shown in the Appx. E.
Table 7: Color and grayscale image denoising PSNR results.
Method # P Color Grayscale
CBSD68 McMaster Urban100 Set12 BSD68 Urban100
@@ 4.2 Evaluation of SemanIR on Various IR Tasks
Evaluation on JPEG CAR. The experiments for color images are conducted with 4 image quality factors ranging from 10 to 40 under two settings (i.e., a single model is trained to handle multiple quality factors, and each model for each quality). The quantitative results shown in Tab. 4 indicate that our SemanIR achieves the best results on all the test sets across various quality factors among all the comparison methods for the color images. The visual comparisons in the Appx. E further supports the effectiveness of our method.
Evaluation on Image Denoising. We show color and grayscale image denoising results in Tab. 7 under two settings (i.e., one model for all noise levels $\sigma=\{15,25,50\}$ and each model for each noise level). For a fair comparison, both parameters and accuracy are reported for all the methods. For , our SemanIR performs better on all test sets for color and grayscale image denoising than others. It is worth noting that we outperform DRUNet and Restormer with lower trainable parameters. For another setting, the proposed SemanIR also archives better results on CBSD68 and Urban100 for color image denoising, and on Set12 and Urban100 for grayscale denoising. These interesting comparisons validate the effectiveness of the proposed SemanIR and also indicate that our method has a higher generalization ability. The visual results in Appx. E also support that the proposed SemanIR can remove heavy noise corruption and preserve high-frequency image details, resulting in sharper edges and more natural textures without over-smoothness or over-sharpness problems.
@@
Evaluation in AWC. We validate SemanIR in adverse weather conditions, including rain+fog (Test1), snow (SnowTest100K), and raindrops (RainDrop). PSNR is reported in Tab. 9. Our method achieves the best performance on Test1 (i.e., 5.76% improvement) and SnowTest100k-L (i.e. 8.01% improvement), while the second-best PSNR on RainDrop compared to all other methods. See Appx. E for Visual comparisons.
Evaluation on Image Demosaicking. The quantitative results shown in 9 indicate that the proposed SemanIR performs best on both the Kodak and MaMaster test sets, especially 0.05dB and 0.45dB absolute improvement compared to the current state-of-the-art.
Evaluation on SR. For the classical image SR, we compared our SemanIR with both recent lightweight and accurate SR models, and the quantitative results are shown in Tab. 10. Compared to EDT, SemanIR-base achieves significant improvements on Urban100 (i.e., 0.72 dB and 0.76dB for 2$\times$ and 4$\times$ SR) and Manga109 datasets (i.e., 0.22dB and 0.17 dB for 2$\times$ and 4$\times$ SR). Even the SemanIR-small consistently ranks as the runner-up across the majority of test datasets, all while maintaining a reduced number of trainable parameters. Visual results in both Fig. 6 and Appx. E also validate the effectiveness of the proposed SemanIR. Specifically, it is clear from the zoomed part in Fig. 6 that SemanIR can restore more details and structural content compared to other methods.